OpenStudy (anonymous):

What is the least common denominator for the equation (4p)/(3p^2-27) = (p+1)/(2p^2+10p+12)

OpenStudy (anonymous):

\[\frac{ 4p }{ 3p ^{2-27} }= \frac{ p +1 }{ 2p ^{2}+10p +12 }\] Is this yoru equation?

OpenStudy (anonymous):

the 27 isn't an exponent.

OpenStudy (anonymous):

or is the minus 27 part of the binomial and not part of the exponent?

OpenStudy (luigi0210):

All you do is factor both out

OpenStudy (anonymous):

part of the binomial

OpenStudy (anonymous):

\[\frac{ 4p }{ 3p ^{2}-27 }=\frac{ p +1 }{ 2p ^{2}+10p +12 }\]

OpenStudy (anonymous):

So first, you factor the denominators of each rational expression.

OpenStudy (anonymous):

Okay. so 3(p^2-9) and ...Idk how to factor trinomials

OpenStudy (anonymous):

\[\frac{ 4p }{ 3\left( p -3 \right)\left( p +3 \right) }=\frac{ p +1 }{2\left( p +2 \right)\left( p +3 \right) }\]

OpenStudy (anonymous):

ok, before that on the left, just factor the 3, you got\[3\left( p ^{2}-9 \right) \] that is called the difference of squares, anythign where you can do a square root of the first term and the second, with a minus sign, factors to (x+y)(x-y) where x is the square root of the first term, and y is the square root of the 2nd term, so x is the square root of x^2, and 3 is the square root of 9

OpenStudy (anonymous):

Ohhh.

OpenStudy (anonymous):

The way to factor a trinomial of the form x^2 +bx +c, you have to find factors of C that add to get bx, so basically \[2p ^{2}+10p +12\] \[2\left( p ^{2}+5p +6 \right)\] Factor out the 2 Now, you need to find 2 factors of 6, so they multiply to get 6, and add up to get 5 Those are "2" and "3" You take those two factors and add them to teh square root of the first term So, \[2\left( p +2 \right)\left( p +3 \right)\]

OpenStudy (anonymous):

With me so far?

OpenStudy (anonymous):

Yeah. Kinda

OpenStudy (anonymous):

There's no real other way to explain how to factor a trinomial. That's just what you have to remember. You need to find 2 numbers that multiply to get the last term, and add up to get the middle term. It can help to make a table for bigger numbers, but for "6", There's only 4 numbers that are divisible 1, 1X6=6 2, 2X3=6 3, 3X2=6 6, 6X1=6 So 1, 2, 3, and 6. Which two also add up to get 5? 3 and 2

OpenStudy (anonymous):

So, now you have \[\frac{ 4p }{ 3\left( p-3 \right)\left( p+3 \right) }=\frac{ p+1 }{ 2\left( p+2 \right)\left( p+3 \right) }\]

OpenStudy (anonymous):

The (p+3) are mentioned twice, so only use one of them You have (p+2) You have (P-3) You have 3 You have 2 OR \[\left( 3 \right)\left( 2 \right)\left( p-3 \right)\left( p+3 \right)\left( P+2 \right)\]

OpenStudy (anonymous):

You can just make the 3 and 2 into a 6, so the LCD is \[6\left( p-3 \right)\left( p+3 \right)\left( p+2 \right)\]

OpenStudy (anonymous):

Do you need to solve the equation too?

OpenStudy (anonymous):

You'll want to practice factoring a whole bunch. Factoring binomials Factoring trinomials Factoring quadranomials by grouping, etc. You want to master factoring before you move onto anything else, because it basically needs to be 2nd nature to you before you move on to stuff like this.

OpenStudy (anonymous):

Okay, thank you